The three non-parallel edges of a rectangular solid are in the ratio 1 : 2 : 3. The diagonal of the solid is 2√14m. Find the length of the longest side.


1
Expert's answer
2021-02-24T06:57:03-0500

let's suppose that the edges of the solid are a, b, c and the diagonal is d. The edges b and c are the sides of the solid base. The diagonal of the base of the solid is e.

Since we have ratio between the edges let's suppose the following:

"a:b:c = 1:2:3;\na = x, b= 2x, c= 3x"


Let's find the diagonal of the base e using the Pythagorean theorem:

"e^2 = a^2 + b^2 ; e^2= 4x^2 + 9x^2"

"e^2 = 13x^2"

Let's use the Pythagorean theorem in the triangle that is build by the edge e, diagonal e and solid diagonal d:

"d^2 = c^2 + e^2"

"(2\u221a14)^2 = x^2 + 13x^2"

"4* 14 = 14x^2"

"4 = x^2; x =2"

The question was to find the longest side:

"a= 2, b= 4, c= 6"


Answer: 6m